OpenStudy (anonymous):

For what values of p does px^2+(2p+1)x+p=0. Have equal roots show all working will give medal

OpenStudy (campbell_st):

well you need to use the discriminant.. which is \[\Delta = b^2 - 4ac\] for equal roots you have the discriminant set to zero so \[b^2 - 4ac = 0\] in your question a = p, b = (2p + 1) and c = p just substitute and then solve for p hope this helps..

OpenStudy (campbell_st):

the equal roots means that the quadratic is a perfect square. and the discriminant is part of the general quadratic formula.